OpenStudy (anonymous):

Please Help! When Θ = 5 pi over 6, what are the reference angle and the sign values for sine, cosine, and tangent?

Nnesha (nnesha):

 these are the reference angles

Nnesha (nnesha):

do you know where is 5pi/6 at in unit circle ?

OpenStudy (anonymous):

Thank you! Here are the multiple choice options by the way

OpenStudy (anonymous):

And I think it's in the second quadrant, right?

Nnesha (nnesha):

yes that's right i like degrees if you like it too then you can convert radians to degrees \[\huge\rm \frac{ 5\pi }{ 6 } \times \frac{ 180 }{ \pi }\]

Nnesha (nnesha):

and if you look at the unit circle 5pi/6 is here and you need signs of cos sin and tan values for that just remember CAST rule or it is like normal x y graph